OFultos The data suggests that the omerence between mean stopping dstances is less than t -10. table in the Appendix of Tables to answer this question r-1292, and i-S59. Caarate·95% a for th, diference between trn overage stopping estances for cas oped with system i and cars eg pped wen system 2 (Round your answers to deow places. 》 -21 70
Answer #1

We have to test (Left tailed T-test)

H_0:\mu _1-\mu _2=-10\\ H_1:\mu _1-\mu _2 < -10

The sample means are \overline{x}=115.2, \overline{y}=129.6 . The sample standard deviations are s s_1=5.06,s _2=5.37 . Sample sizes are  n_1=7,n_2=7
The estimated mean difference is \overline{x}-\overline{y}=-14.4
We use T distribution since the sample sizes are small. The test statistic

t=\left (\overline{x}-\overline{y} +10\right )/\sqrt{\frac{s _1^2}{n_1}+\frac{s _2^2}{n_2}}\\ {\color{Blue} t=-1.58}

The statistic follows T-distribution.

The P-value is P\left ( T<t \right )=P\left (T<-1.58\right )={\color{Blue} 0.070}

Sinc the P-value \textup{P-value} = 0.07>\alpha =0.01   we retain retain the null hypothesis

Fail to reject the null hypothesis.

We are required to solve only one question. Please post the remaining questions as another post.